OpenStudy (bethgrace96):

MEDAL WILL BE REWARDED:A ball is thrown upward. its height ( h, in feet ) is modeled by the function h = -16t^2 + 64t+3, where t is the length of time ( in seconds ) that the ball has been in the air. what is the maximum height the ball reaches?

OpenStudy (tkhunny):

Have you considered \(\dfrac{-64}{2(-16)}\)? aka -b/(2a) Have you considered Completing the Square?

OpenStudy (bethgrace96):

no

OpenStudy (tkhunny):

Oh, well, why don't you do that!?

OpenStudy (bethgrace96):

the answer i got was 2 and thats not an answer choice. answer choices are : 3, 51,63,67

ganeshie8 (ganeshie8):

Yes, the maximum value occurs at \(t = 2\)

ganeshie8 (ganeshie8):

to get the maximum height, evaluate the given height function at \(t = 2\)

OpenStudy (bethgrace96):

how?

ganeshie8 (ganeshie8):

h = -16t^2 + 64t+3

ganeshie8 (ganeshie8):

plugin t = 2 above^

OpenStudy (bethgrace96):

okay one second

OpenStudy (bethgrace96):

i dont understand.

OpenStudy (bethgrace96):

where do i plug the 2 in

ganeshie8 (ganeshie8):

h = -16t^2 + 64t+3 plugin t = 2, you get : h = -16*2^2 + 64*2 +3

ganeshie8 (ganeshie8):

simplify

OpenStudy (bethgrace96):

i did. its a huge number

OpenStudy (bethgrace96):

i got 643?

ganeshie8 (ganeshie8):

you should get 67

OpenStudy (bethgrace96):

what did i do wrong

ganeshie8 (ganeshie8):

not sure, check that wolfram link...